e^x dx+(e^x coty +2ycscy)dy=0
1
Expert's answer
2020-10-05T16:53:50-0400

The given differential equation is

"e^xdx+(e^xcoty+2ycosecy)dy = 0."


Now if we multiply both sides by "siny , we \space get"


"e^xsiny \space dx + (e^xcoty *siny + 2y *cosecy*siny)\space dy =0\\or, e^xsiny \space dx + (e^xcosy + 2y) \space dy=0\\or, e^x siny \space dx +e^xcosy \space dy+ 2ydy = 0\\"


Now this is much more simplified form of the given differential equation and can be written as

"d(e^xsiny)+ d(y^2) = 0"

Now integrating botth sides we get,

"\int d(e^xsiny) + \int d(y^2) = C" ( where C is a constant of integration)


"e^xsiny + y^2 = C"


"\therefore" The required solution is

"e^xsiny + y^2 = C"